# Causal Learning for Trustworthy Recommender Systems: A Survey

### v1

@@ Abstract
Recommender Systems (RS) have significantly advanced online content discovery and personalized decision-making. However, emerging vulnerabilities in RS have catalyzed a paradigm shift towards Trustworthy RS (TRS). Despite numerous progress on TRS, most of them focus on data correlations while overlooking the fundamental causal nature in recommendation. This drawback hinders TRS from identifying the cause in addressing trustworthiness issues, leading to limited fairness, robustness, and explainability. To bridge this gap, causal learning emerges as a class of promising methods to augment TRS. These methods, grounded in reliable causality, excel in mitigating various biases and noises while offering insightful explanations for TRS. However, there lacks a timely survey in this vibrant area. This paper creates an overview of TRS from the perspective of causal learning. We begin by presenting the advantages and common procedures of Causality-oriented TRS (CTRS). Then, we identify potential trustworthiness challenges at each stage and link them to viable causal solutions, followed by a classification of CTRS methods. Finally, we discuss several future directions for advancing this field.
@@ 1 Introduction
Recommender Systems (RS), as one of the most successful applications of artificial intelligence, have revolutionized the way of our decision-making in the digital world. However, with their growing influence, concerns regarding the trustworthiness of RS have become increasingly prevalent. Inspecting the lifecycle of RS Wang et al. (2023), we identify several significant trustworthiness challenges across different stages, especially regarding fairness, robustness and explainability. They present pivotal yet challenging objectives for modern RS, extending beyond accuracy. Specifically, fairness requires unbiased decision-making for various users, items and providers. Robustness emphasizes the reliability of RS when confronted with noises or attacks. Explainability poses an urgent need for strong interpretability, advancing the understanding of RS mechanisms and results. Recently, due to its exceptional ability in identifying and utilizing causality, going beyond correlations, causal learning has been considered an effective and elegant tool to analyze and address Trustworthy RS (TRS) problems Gao et al. (2024); Li et al. (2023b). In this paper, we overview the trustworthiness challenges throughout the RS lifecycle, alongside a comprehensive exploration of causality-oriented solutions.
Figure 1: Illustrative comparison between conventional RS and CTRS. CTRS enhances robustness by learning underlying causality, whereas conventional RS may be misled by unreliable correlations.
@@ 1.0.1 Motivation: why Causality-oriented TRS (CTRS)?
Causality is more reliable than correlations. Conventional RS heavily rely on correlations learned from observed data. However, correlations are not always causally correct and spurious correlations can be notably misleading, as exemplified in Figure 1. In contrast, CTRS methods He et al. (2023) are adept at identifying and modeling genuine causal relations for enhanced robustness. Moreover, another persistent issue in RS is the non-random nature of missing data Wang et al. (2019), as user-item interactions are typically preference-driven, rather than random. By leveraging advanced causal techniques (e.g., propensity reweighting Li et al. (2023a)), the impacts of this issue can be effectively mitigated, paving the way for more accurate and reliable recommendations.
@@
Causal learning helps provide straightforward and effective explanation. Conventional RS models, which are commonly considered black-box models, can hardly offer clues about the reason for recommendations. CTRS, on the other hand, empowered by path discovery and counterfactual inference, possess superior interpretability. These methods Xian et al. (2019); Yu et al. (2023) not only shed light on the recommendation mechanism but also indicate the rationale behind the results, thereby providing more transparent and comprehensible frameworks for both users and developers.
@@ 1.0.2 How does causal learning help solve TRS problems?
Causal learning shows promising efficacy in understanding the cause-and-effect relationships of variables in a TRS problem. Typical CTRS methods include two major steps, i.e., causal formulation and causal inference. The former step aims to accurately delineate causal relationships within the system. This is usually achieved by the Structural Causal Model (SCM) Pearl (2009). Formally, it is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is a set of nodes, representing variables associated with the problem, and $\mathcal{E}$ is a set of edges, illustrating underlying causal relationships. Following such established causal models, either through the process of causal discovery or leveraging prior knowledge, the root causes of problems related to trustworthiness can be clearly identified. Subsequently, advanced causal inference tools can be applied to intervene in causality and mitigate effects of these problems.
Despite several surveys on TRS Wang et al. (2023); Ge et al. (2022); Fan et al. (2022), as well as causal learning related RS Gao et al. (2022); Luo et al. (2023); Zhu et al. (2023); Xu et al. (2023b); Saito and Joachims (2021); Wu et al. (2022b), they lack a specific focus on CTRS. Surveys on TRS typically concentrate on data-driven methods while these methods are prone to issues of spurious correlations, biases and limited interpretability, as discussed earlier. Meanwhile, surveys on causal learning for RS often focus on technical causal inference tools applied in RS, neglecting the connection with trustworthiness challenges. Moreover, most of them overlook the importance of ensuring trustworthiness across different stages. To fill these gaps, we provide a structured survey to examine trustworthiness challenges throughout the RS lifecycle and align them with causal solutions.
Contributions. The main contributions of this survey are summarized as follows:
  • It outlines trustworthiness challenges in the lifecycle of RS and aligns them with viable causal solutions, highlighting the significance and necessity of causal learning in addressing TRS problems.
  • It summarizes the current progress of CTRS, aiming to hold practical values in guiding the development of unified trustworthy RS platforms.
  • It offers valuable insights into future directions and perspectives for the future advancement of CTRS.
RS stage
Aspect
Sub-aspect
Trustworthiness challenge
Causal solution
Example
Data preparation
Fairness
User side
User attribute bias
Counterfactual inference
[1]
Population bias
Counterfactual augmentation
Item side
Item attribute bias
Counterfactual inference
[1]
Popularity bias
Structure adjustment
[2], [3]
Counterfactual inference
[4]
Position bias
Propensity reweighting
[5]
Robustness
Non-randomness
Selection bias
Propensity reweighting
[6], [7], [8]
Corruption
Poisoning attacks
Counterfactual augmentation
Representation learning
Fairness
Implicit intention
Conformity bias
Causal disentanglement
[9]
Robustness
Noise
Spurious correlations
Structure adjustment
[10]
Counterfactual augmentation
[11], [12]
Noisy feedback
Counterfactual inference
[13]
Recommendation generation
Fairness
Model side
Bias amplification
Structure adjustment
[14]
Filter bubble bias
Counterfactual inference
[15], [16]
Explainability
Model agnostic
Unexplainable results
Counterfactual inference
[17], [18], [19]
Model specific
Black-box mechanism
Path discovery
[20]
Evaluation
Robustness
Model effect
Uplift evaluation
Uplift effect estimation
[21]
[1]Li et al. (2021); [2]Zhang et al. (2021c); [3]Gupta et al. (2021); [4]Wei et al. (2021); [5]Joachims et al. (2018); [6]Zhang et al. (2020b); [7]Li et al. (2023b); [8]Li et al. (2023a); [9]Zheng et al. (2021); [10]He et al. (2023); [11]Mu et al. (2022); [12]Xu et al. (2023a); [13]Zhang et al. (2021b); [14]Wang et al. (2021b); [15]Wang et al. (2022b); [16]Gao et al. (2024); [17]Tan et al. (2021); [18]Chen et al. (2022); [19]Yu et al. (2023); [20]Xian et al. (2019); [21]Sato et al. (2019)
@@ 2 Trustworthiness Challenges in RS Lifecycle
A typical recommender system is mainly composed of four successive and inter-connected stages, i.e., data preparation, representation learning, recommendation generation and evaluation. Each stage plays a fundamental role in the lifecycle of RS Wang et al. (2023) and is confronted with its own trustworthiness challenges. We highlight that the overall trustworthiness of RS essentially hinges on the reliability of each stage. Only when each stage is deemed trustworthy can we develop a unified trustworthy RS. In this section, we delve into the trustworthiness challenges inherent at each stage, examining them through the lens of causal learning. Table 1 summarizes these challenges across the various stages and illustrates their alignment with causal solutions, offering a structured overview of the landscape.
@@ 2.1 Challenges in Data Preparation
As the first stage of RS, data preparation involves gathering, organizing and transforming input data (e.g., user-item interactions and side information). However, raw data collected from complex cyberspace can be insufficiently trustworthy and commonly carry various biases (e.g., popularity bias), noises (e.g., fake ratings/reviews from online water army) and even malicious attacks. This results in substantial challenges to the overall trustworthiness of the system, especially in ensuring fairness and robustness in data preparation.
@@ 2.1.1 Fairness Challenges in Data Preparation
During data preparation, a primary fairness challenge lies in how to mitigate the pre-existing biases Li et al. (2023c) in raw data. Neglecting such inherent biases will pose serious fairness risks to those RS models trained on these biased data, potentially harming the interests of stakeholders. According to different subjects of fairness concerns, there exist two major types of pre-existing biases, i.e., user-side Li et al. (2021) and item-side bias Zhang et al. (2021c).
@@
User-side bias refers to unfair recommendations for different groups of users. The first prevalent bias in this category, i.e., user attribute bias, refers to unequal treatment of users based on their sensitive attributes (e.g., gender, race, and educational background). For example, a biased job recommender may make discriminatory decisions based on the gender of users Mansoury et al. (2020a). Another critical user-side bias, namely population bias, stems from the unbalanced distribution of users. Typically, heavy users contribute a significantly larger volume of interaction data than casual users, leading to a highly unbalanced dataset. Consequently, RS models trained on such data tend to favor heavy users and perform noticeably better Wang et al. (2022a); Wu et al. (2022a) for them than for casual users.
@@
Additionally, in a trustworthy recommendation ecosystem Wang et al. (2023), mitigating item-side bias helps to safeguard the interests of items and their providers. Similar to users, items may suffer from attribute bias because of certain properties. However, beyond general attributes, such as brands and production places, the popularity of an item can markedly affect the recommendation results, giving rise to popularity bias Zhang et al. (2021c). Popular items, benefiting from their fame, naturally draw more attention than long-tail products. If RS models capture and rely on these patterns, they may blindly prefer to recommend popular items without adequately considering the user-item alignment. Another classic item-side bias, namely position bias Joachims et al. (2018), is based on the common observation Joachims et al. (2007) that users tend to pay more attention and are more likely to interact with items at the top of a recommendation list, regardless of their actual relevance or quality. Therefore, such biased interaction data may mislead newly trained models, resulting in unfairness to other items.
Several efforts Wu et al. (2021) have been made in recent years to address the aforementioned pre-existing biases. For instance, Abdollahpouri et al. Abdollahpouri et al. (2017) developed a regularization-based framework to achieve a balance between accuracy and fairness objectives. However, these data-driven methods, relying on patterns observed in data, often fall short in understanding the mechanism of how bias affects interactions. This significantly limits their efficacy in debiasing within practical applications. To bridge this gap, several causality-oriented methods have been proposed, focusing on analyzing and measuring the causal relationships within a biased RS. For user-side bias, counterfactual reasoning stands out as an elegant solution, which can not only achieve counterfactual fairness Li et al. (2021) for individuals by reducing the impacts of sensitive attributes, but also can help to generate reasonable samples Mu et al. (2022) for minority users. For popularity bias, existing work commonly formulates the problem within the framework of SCM and performs structure adjustment Zhang et al. (2021c) or counterfactual inference Wei et al. (2021) for debiasing. For position bias, Joachims et al. Joachims et al. (2018) devised a propensity-weighted model to reduce the bias caused by positions in ranking. The details of causal learning methods will be introduced in Section 3.
@@ 2.1.2 Robustness Challenges in Data Preparation
Due to the inherent high sparsity of user-item interactions, most existing work assumes that interactions are Missing At Random (MAR) for modeling simplicity and statistical convenience. However, this assumption does not always hold in practice. In cases where samples are Missing Not At Random (MNAR), selection bias Li et al. (2023b) becomes a significant issue. It refers to the distortion that data used to train RS models are not representative of the entire range of user preferences. To overcome this challenge, earlier work Hernández-Lobato et al. (2014) managed to model the generative and missing process of data, but this led to high computational costs. Recently, propensity reweighting tools have been widely adopted to solve selection bias in RS. These methods offer higher efficiency Schnabel et al. (2016) and flexibility to integrate with various recommendation frameworks Zhang et al. (2020b); Li et al. (2023a).
Another critical robustness challenge in data preparation is how to defend against poisoning attacks Zhang et al. (2022). These attacks involve the insertion of fake users into training data, with the intent to manipulate the outcomes of RS models for malicious purposes. To counteract such threats, two representative defense strategies are primarily adopted, i.e., detection-based methods Zhang et al. (2020a) and adversarial training Tang et al. (2020). Detection-based methods focus on developing an accurate classifier to identify anomalies and spurious samples. Adversarial training, on the other hand, aims to enhance model robustness by introducing generated perturbations during training. However, these defense methods are essentially data-driven. There remains an imperative need to explore causality-oriented defense methods, pursuing a deeper understanding of how attacks manipulate RS results. Useful causal learning tools include counterfactual augmentation for adversarial training and causal effect inference for causal measurement.

### v2

@@ Abstract
Recommender Systems (RS) have significantly advanced online content filtering and personalized decision-making. However, emerging vulnerabilities in RS have catalyzed a paradigm shift towards Trustworthy RS (TRS). Despite substantial progress on TRS, most efforts focus on data correlations while overlooking the fundamental causal nature of recommendations. This drawback hinders TRS from identifying the root cause of trustworthiness issues, leading to limited fairness, robustness, and explainability. To bridge this gap, causal learning emerges as a class of promising methods to augment TRS. These methods, grounded in reliable causality, excel in mitigating various biases and noise while offering insightful explanations for TRS. However, there is a lack of timely and dedicated surveys in this vibrant area. This paper creates an overview of TRS from the perspective of causal learning. We begin by presenting the advantages and common procedures of Causality-oriented TRS (CTRS). Then, we identify potential trustworthiness challenges at each stage and link them to viable causal solutions, followed by a classification of CTRS methods. Finally, we discuss several future directions for advancing this field.
@@ 1 Introduction
Recommender Systems (RS), as one of the most successful applications of artificial intelligence, have revolutionized the way of our decision-making in the digital world. However, with their growing influence, concerns over trustworthiness of RS have become increasingly prevalent. Inspecting the lifecycle of RS, i.e., data preparation, representation learning, recommendation generation, and evaluation, Wang et al. (2023a), we identify several significant trustworthiness challenges, which are the obstacles that hinder the system’s fairness, robustness and explainability. They present pivotal yet challenging objectives for modern RS, extending beyond accuracy. Specifically, fairness requires unbiased decision-making for various users, items, and providers. Robustness emphasizes the reliability of RS against noise or attacks. Explainability poses an urgent need for strong interpretability, advancing the understanding of RS mechanisms and results. Recently, causal learning has emerged as a powerful tool for addressing Trustworthy RS (TRS) challenges Gao et al. (2024b); Li et al. (2023b), leveraging causality rather than mere correlations. In this paper, we overview the trustworthiness challenges throughout the RS lifecycle, alongside a comprehensive investigation of causality-oriented solutions.
Figure 1: Conventional RS relies on correlations, often misled by popularity bias (e.g., suggesting trending thrillers to users who prefer period dramas). CTRS performs causal interventions to reduce bias, ensuring more relevant and trustworthy recommendations.
@@ 1.0.1 Motivation: why Causality-oriented TRS (CTRS)?
Causality is more reliable than correlations. Conventional RS heavily relies on correlations learned from observed data. However, correlations are not always causally correct, and spurious correlations can be notably misleading, as exemplified in Figure 1. In contrast, CTRS methods He et al. (2023) are adept at identifying genuine causal relations and applying causal adjustments for enhanced trustworthiness. Moreover, another persistent issue in RS is the non-random nature of missing data Li et al. (2023b), as user-item interactions are typically preference-driven. Causal techniques like propensity reweighting Li et al. (2023a) can help mitigate this issue, leading to more accurate and reliable recommendations.
@@
Causal learning helps provide a straightforward and effective explanation. RS models, especially deep learning-based ones, often lack interpretability, and thus can hardly offer clues about the reason for recommendations. In contrast, CTRS leverages path discovery and counterfactual inference for superior interpretability. These methods Xian et al. (2019); Wang et al. (2024a) not only illuminate recommendation mechanisms but also clarify the rationale behind recommendation results, thereby providing more transparent and comprehensible frameworks for both users and developers.
@@ 1.0.2 How does causal learning help solve TRS problems?
Causal learning enhances TRS by explicitly modeling, analyzing, and intervening in the underlying cause-and-effect relationships. Specifically, it operates through two key steps: causal formulation and causal inference. The first step delineates causal relationships within the system by constructing causal models Rubin (1974), such as Structural Causal Models (SCM) Pearl (2009). It identifies relevant variables (e.g., user preferences, item attributes) and maps their causal influences, either through causal discovery or domain expertise. In the second step, causal techniques are applied to estimate and adjust for causal effects. Thus, we can move beyond correlation-based models to understand and manipulate the causal factors affecting trustworthiness for robust, fair, and explainable recommendations.
1.0.3 Comparison with Existing Surveys
Despite several surveys on relevant topics, three critical gaps remain: 1) Limited scope: TRS surveys Fan et al. (2022); Ge et al. (2024); Wang et al. (2023a) primarily review general aspects of trustworthiness in RS, such as fairness, robustness, and explainability. However, they lack a particular focus on causality-oriented TRS, which is a significant branch of TRS and cannot be overlooked. 2) Coarse-grained challenge analysis: Existing literature Fan et al. (2022); Li et al. (2023c) typically examines TRS challenges at a high level from general aspects, without mapping them to specific stages of RS lifecycle. Hence, it lacks insights into building a practical and unified TRS platform, as trustworthiness concerns vary across different stages in real-world applications. 3) Lack of causal alignment: While surveys on causal inference based RS Gao et al. (2024a); Luo et al. (2023); Zhu et al. (2023); Xu et al. (2023a); Wu et al. (2022) focus on technical methodologies, they fail to explicitly align causal learning with trustworthiness challenges or provide insights on why and how causal learning can help build TRS.
Contributions. To bridge these gaps, we provide a dedicated survey on CTRS to examine trustworthiness challenges and corresponding causal solutions throughout the RS lifecycle. The main contributions are summarized as follows:
  • Granular, stage-wise challenge analysis: We introduce an in-depth examination and a novel taxonomy of key challenges in building TRS. Through well mapping different aspects of challenges to different stages when building RS, we offer a structured overview and actionable insights for addressing specific trustworthiness issues at each RS stage.
  • Comprehensive CTRS method summary and comparison: We provide an in-depth survey on CTRS methods, highlighting the unique advantages of causal learning in tackling TRS challenges. In addition, a critical comparison of CTRS methods is provided to guide the selection of appropriate solutions for addressing specific challenges.
  • Insightful challenge-method alignment: We highlight the underlying connections and alignment between trustworthiness challenges and causal learning methods, explaining why and how different causal learning methods address corresponding challenges. In addition, we offer valuable perspectives on future directions for advancing CTRS.
RS stage Aspect Sub-aspect Trustworthiness challenge Causal solution Example
Data preparation Fairness User side User attribute bias Counterfactual inference [1], [2]
Population bias Counterfactual augmentation -
Item side Item attribute bias Counterfactual inference [1], [2]
Popularity bias Structure adjustment [3], [4]
Counterfactual inference [5]
Position bias Propensity reweighting [6]
Robustness Non-randomness Selection bias Propensity reweighting [7], [8], [9]
Corruption Poisoning attacks Counterfactual augmentation -
Representation learning Fairness Implicit intention Conformity bias Causal disentanglement [10]
Robustness Noise Spurious correlations Structure adjustment [11]
Counterfactual augmentation [12], [13]
Noisy feedback Counterfactual inference [14]
Recommendation generation Fairness Model side Bias amplification Structure adjustment [15]
Filter bubble bias Counterfactual inference [16], [17]
Explainability Model agnostic Unexplainable results Counterfactual inference [18], [19], [20]
Model specific Black-box mechanism Path discovery [21]
Evaluation Robustness Model effect Uplift evaluation Uplift effect estimation [22]
[1]Li et al. (2021); [2]Chen et al. (2024); [3]Zhang et al. (2021c); [4]Gupta et al. (2021); [5]Wei et al. (2021); [6]Joachims et al. (2018); [7]Liu et al. (2023); [8]Li et al. (2023b); [9]Li et al. (2023a); [10]Zheng et al. (2021); [11]He et al. (2023); [12]Mu et al. (2022); [13]Xu and et al. (2023); [14]Zhang et al. (2021b); [15]Wang et al. (2021b); [16]Wang et al. (2022b); [17]Gao et al. (2024b); [18]Tan et al. (2021); [19]Wang et al. (2024a); [20]Wang et al. (2024b); [21]Xian et al. (2019); [22]Sato et al. (2019)
@@ 2 Trustworthiness Challenges in RS Lifecycle
2.1 Problem Formulation
Recommender systems aim to predict user preferences over items to provide personalized recommendations. Formally, let $\mathcal{U}=\{u_{1},u_{2},\cdots,u_{M}\}$ denote a set of $M$ users, and =$\mathcal{I}=\{i_{1},i_{2},\cdots,i_{N}\}$ denote a set of $N$ items. The observed interactions between users and items are represented as $\mathcal{D}=\{(u,i,r_{ui})\mid u\in\mathcal{U},i\in\mathcal{I}\}$, where $r_{ui}$ is the explicit rating or implicit feedback from user $u$ on item $i$. The goal of an RS is to learn a predictive function $f$, parameterized by $\Theta$, with user and item representations $\mathbf{h}_{u}$ and $\mathbf{h}_{i}$, such that $\hat{r}_{ui}=f(\mathbf{h}_{u},\mathbf{h}_{i}\mid\Theta)$ estimates the true preference.
To achieve this goal, a typical RS comprises four successive and inter-connected stages, i.e., data preparation, representation learning, recommendation generation, and evaluation. Each stage plays a fundamental role in the RS lifecycle Wang et al. (2023a) and is confronted with its own trustworthiness challenges. We highlight that the overall trustworthiness of RS essentially hinges on the reliability of each stage. Only when each stage is deemed trustworthy can we develop a unified trustworthy RS. In this section, we delve into the trustworthiness challenges inherent at each stage, examining them through the lens of causal learning. Table 1 summarizes these challenges across the various stages and illustrates their alignment with causal solutions, offering a structured overview of the landscape.
@@ 2.2 Challenges in Data Preparation
As the first stage of an RS, data preparation involves gathering, organizing and transforming input data $\mathcal{D}$ (e.g., user-item interactions and side information). However, raw data collected from complex cyberspace often contain various biases (e.g., popularity bias), noise (e.g., fake ratings/reviews from online water armies), and even malicious attacks, posing significant risks to the system’s fairness and robustness.
@@ 2.2.1 Fairness Challenges in Data Preparation
During data preparation, a primary fairness challenge lies in how to mitigate the pre-existing biases Li et al. (2023c) in raw data $\mathcal{D}$. Such inherent biases directly affect RS models trained on these data, potentially harming the interests of stakeholders. According to different subjects of fairness concerns, there exist two major types of pre-existing biases, i.e., user-side Li et al. (2021) and item-side bias Li et al. (2025).
User-side bias refers to unfair recommendations for different groups of users. The first prevalent bias in this category is known as user attribute bias. Let $A_{u}$ denote a sensitive attribute associated with user $u$, such as gender, age, or ethnicity. User attribute bias occurs when RS models make discriminatory decisions based on these sensitive attributes, i.e.,
$\mathbb{E}[\hat{r}_{ui}\mid A_{u}=a]\neq\mathbb{E}[\hat{r}_{ui}\mid A_{u}=b],$ (1)
@@
for different values $a$ and $b$ of $A_{u}$. For example, a biased job recommender may make discriminatory decisions based on the gender of users Mansoury et al. (2020a). Another critical user-side bias, namely population bias, stems from the unbalanced distribution of user data. Define subsets of users based on their activity levels, we have heavy users $\mathcal{U}_{\rm heavy}$ who are frequently active and casual users $\mathcal{U}_{\rm casual}$ who interact sporadically. Observe that $\mathcal{U}_{\rm heavy}$ typically contributes a significantly larger volume of interaction data than $\mathcal{U}_{\rm casual}$, i.e., $|\mathcal{D}_{\rm heavy}|\gg|\mathcal{D}_{\rm casual}|$, where $\mathcal{D}_{\rm heavy}=\{(u,i,r_{ui})\mid u\in\mathcal{U}_{\rm heavy}\}$ and $\mathcal{D}_{\rm casual}=\mathcal{D}\setminus\mathcal{D}_{\rm heavy}$. Thus, RS models trained on such data tend to favor heavy users and perform noticeably better for them than for casual users Wang et al. (2022a).
@@
Apart from user-side bias, in a trustworthy recommendation ecosystem Wang et al. (2023a), mitigating item-side bias helps to safeguard the interests of items and their providers. Similar to users, items may suffer from attribute bias due to certain attributes $A_{i}$ (e.g., brands and production places). However, beyond general attributes, the popularity of an item can markedly affect the recommendation results, giving rise to popularity bias Klimashevskaia et al. (2024). Popular items $\mathcal{I}_{\rm pop}$, benefiting from their fame, naturally draw more attention than long-tail ones $\mathcal{I}_{\rm long-tail}$. The bias is shown as:
$\mathbb{E}[\hat{r}_{ui}\mid i\in\mathcal{I}_{\rm pop}]>\mathbb{E}[\hat{r}_{ui}\mid i\in\mathcal{I}_{\rm long-tail}],$ (2)
even when users do not necessarily prefer popular items. Another classic item-side bias, namely position bias Joachims et al. (2018), is based on the common observation Joachims et al. (2007) that users tend to pay more attention and are more likely to interact with items at the top of a recommendation list, regardless of their actual relevance or quality. Therefore, such biased interaction data may mislead newly trained models, resulting in unfairness to other items.
Traditional methods mitigate these pre-existing biases by resampling or imposing additional constraints during training Wu et al. (2021). However, without understanding the underlying causal mechanisms through how biased data affect the recommendation process, these methods may achieve only superficial fairness and exhibit limited generalization ability. To address this, causality-oriented methods have been proposed, analyzing and measuring the causal relationships within a biased RS. For user-side bias, counterfactual reasoning offers an elegant solution, ensuring counterfactual fairness Li et al. (2021) for individuals with sensitive attributes while generating reasonable samples Mu et al. (2022) for minority users. For item-side bias, existing work often formulates SCM for thorough analysis and mitigates the negative effects through causal interventions Zhang et al. (2021c); Wei et al. (2021). The details of causal learning methods are introduced in Section 3.
@@ 2.2.2 Robustness Challenges in Data Preparation
Due to the inherent high sparsity of user-item interactions, most existing work assumes that interactions are Missing At Random (MAR) for modeling simplicity and statistical convenience. Let $\mathbf{O}\in\{0,1\}^{M\times N}$ be an indicator matrix where each entry $o_{ui}=1$ if $r_{ui}$ is observed, and $o_{ui}=0$ otherwise. The MAR assumption implies that the probability of observing an interaction is independent of the unobserved rating $r_{ui}$. However, this assumption does not always hold in practice. In cases where samples are Missing Not At Random (MNAR), i.e., the missingness depends on the unobserved ratings:
$P(o_{ui}=1\mid r_{ui},\mathbf{X}_{ui})\neq P(o_{ui}=1\mid\mathbf{X}_{ui}),$ (3)
where $\mathbf{X}_{ui}$ represents observed features related to user $u$ and item $i$, selection bias Li et al. (2023b) becomes a significant issue. It refers to the distortion that data used to train RS models are not representative of the entire range of user preferences. To overcome this challenge, earlier work Hernández-Lobato et al. (2014) managed to model the generative and missing process of data, but this led to high computational costs. Recently, propensity reweighting tools have been widely adopted to solve selection bias in RS. These methods offer higher efficiency and flexibility to integrate with other recommendation frameworks Li et al. (2023a).
Another critical robustness challenge at this stage is defending against poisoning attacks Zhang et al. (2022). These attacks usually insert fake users into training data to manipulate RS outcomes for malicious purposes. To counteract such threats, two representative defense strategies are primarily adopted, i.e., detection-based methods Zhang et al. (2020a) and adversarial training Tang et al. (2020). Detection-based methods focus on developing an accurate classifier to identify and filter out spurious samples. Adversarial training, on the other hand, aims to enhance model robustness by introducing generated perturbations during training. However, the generalization ability of these defense methods is constrained by the given training data, whereas examining how attacks manipulate RS results from a causal perspective can help improve it. Useful causal techniques include counterfactual augmentation for adversarial training and causal effect inference for causality measurement.